Write a system of two equations in two unknowns. Solve the system by using the addition method. « Mike paid $2.98 for three doughnuts and two coffees. The next day, he paid $1.64 for two doughnuts and one coffee. How much does it cost for one doughnut and one coffee? »

Answer by checkley77(12844)

3D+2C=2.98
2D+C=1.64 MULTIPLY BY -2 & ADD.
-4D-2C=-3.28
3D+2C=2.98
-----------------
-D=-.30
D=.30 COST OF EACH DOUGHNUT.
3*.30+2C=2.98
.90+2C=2.98
2C=2.98-.90
2C=2.08
C=2.08/2
C=1.04 COST OF EACH COFFEE.
PROOF:
2*.30+1.04=1.64
.60+1.04=1.64
1.64=1.64